### v2

# Rate-distortion Theory with Lower Semi-continuous Distortion: A Concentration-compactness Approach

Abstract
In this paper, we study rate-distortion theory for general sources with an emphasis on the existence of optimal reconstruction distributions. Classical existence results rely on compactness assumptions with continuous distortion that are often violated in general settings. By introducing the concentration-compactness principle into the analysis of the rate-distortion functional, we establish the existence of optimal reconstructions under mild coercivity and lower semi-continuity conditions on the distortion function. Our results provide a unified and transparent existence theorem for rate-distortion problems with lower semi-continuous distortion.
I Introduction
Rate-distortion (RD) theory is a fundamental branch of information theory, primarily concerned with the problem of compressing data at the minimum coding rate subject to a given distortion constraint. Initially proposed by Claude E. Shannon in his groundbreaking works in 1948 [35] and 1959 [34], this theory has since become a cornerstone of modern data compression and communication system design. RD problems focus on designing a coding scheme that minimizes the coding rate, while satisfying a specific distortion level for a given source and distortion measure.
Classical formulations of the rate-distortion problem primarily focus on finite-alphabet sources or continuous sources defined on Euclidean spaces. Comprehensive treatments can be found in the monographs of Berger [2], Blahut [4, 3], and Cover and Thomas [10]. A major conceptual advance was achieved by I. Csiszár, who extended RD theory to abstract alphabets [11]. He rigorously derived the parametric representation of the RD function, laying a solid theoretical foundation for numerical algorithms [4, 1], and also established the existence of optimal reconstruction distributions under the assumption that the reproduction space is compact. Further contributions were made by E. Riegler et al. [32, 33], who focused on lossy compression theory for sources defined on compact manifolds and fractal sets. By leveraging results related to the rate-distortion dimension [20], they derived the Shannon lower bound (SLB) in more general settings, thereby extending the applicability of RD theory to more complex source structures. In terms of the parametric representation of the RD function, V. Kostina and E. Tuncel [21] avoided reliance on the Karush-Kuhn-Tucker (KKT) optimality conditions. Instead, they streamlined I. Csiszár’s argument by utilizing the Donsker-Varadhan characterization of minimum relative entropy [14]. Their work significantly advanced the theory of successive refinement for abstract sources, thereby generalizing the earlier results of Equitz and Cover [15]. More recently, connections between rate-distortion theory and optimal transport have attracted increasing attention. In particular, the framework of optimal weak transport (OWT) [17] provides a natural formulation for problems in which the optimization is performed over disintegrations of measures rather than fixed couplings. This perspective has been shown to recover several structural properties of the rate-distortion function and offers a complementary geometric viewpoint on lossy source coding [38].
Despite these developments, the question of existence of optimal reconstructions in general non-compact settings remains delicate. In many applications of practical and theoretical interest, the reproduction alphabet is naturally modeled as a non-compact Polish space, such as $\mathbb{R}^{n}$ or an infinite-dimensional function space. In such settings, classical compactness-based arguments, most notably those relying on Prokhorov’s theorem, no longer apply directly. As a consequence, minimizing sequences of the rate-distortion functional may fail to admit convergent subsequences, exhibiting pathological behaviors such as mass escaping to infinity or splitting into widely separated components.
A significant step toward relaxing compactness assumptions was made by Rezaei et al. [31], who employed Stone-Čech compactification and fixed-point arguments in locally convex topological vector spaces to establish existence results under weaker topological conditions. However, their approach requires continuity assumptions on the distortion function. In contrast, in general rate-distortion problems one typically assumes only measurability of the distortion function (in a standard Borel setting, which is necessary for the relative entropy to be well-defined; see Example 2.5 in [29] for details). This gap motivates the present work, where we investigate the existence of optimal reconstructions under substantially weaker regularity assumptions on the distortion function.
In this paper, we address this existence problem from a variational and measure-theoretic perspective. Our central observation is that the fundamental obstruction arises from a loss of compactness in the space of reconstruction measures when the reproduction alphabet is non-compact. Such loss of compactness manifests itself through possible dispersion phenomena of minimizing sequences, which cannot be controlled by standard tightness arguments. To overcome this difficulty, we introduce the concentration-compactness principle developed by Lions [23, 24, 25, 26], a powerful tool in nonlinear functional analysis designed precisely to handle variational problems where standard compactness fails. By applying concentration-compactness to the rate-distortion functional, we obtain a precise characterization of all possible asymptotic behaviors of minimizing sequences, corresponding to compactness (or tightness), vanishing, or dichotomy. Under natural and verifiable coercivity conditions on the distortion function, we rule out the latter two scenarios and thereby establish the existence of an optimal reconstruction distribution without assuming compactness of the reproduction space. This leads to a more direct and transparent existence theorem for rate-distortion minimizers on general proper Polish spaces.
The remainder of this paper is organized as follows. Section II reviews the necessary background on rate-distortion theory, including classical existence results and variational representations. In Section III, we present our main existence theorems for optimal reconstruction distributions on proper Polish spaces, together with a generalized concentration-compactness principle and suitable coercivity conditions on the distortion function. The proofs of the main results are provided in the appendices, where we give a detailed analysis of minimizing sequences via concentration-compactness arguments. In Section IV, we conclude with a discussion of potential extensions and applications of our framework.
II A Recap of Rate-Distortion Theory
Rate-distortion theory is a central branch of information theory, concerned with characterizing the fundamental limits of source compression when a prescribed level of distortion is permitted. Given a source random variable $X$ with probability distribution $p(x)$ and a distortion (or loss) function $\rho(x,y)$, the rate-distortion function $R(D)$ is defined as the minimum information rate required to represent the source subject to an average distortion constraint $D$:
$R(D)=\inf_{p(y|x)}I(X;Y),$ (1)
where the infimum is taken over all conditional distributions (or disintegrations) satisfying the fidelity constraint
$\mathbb{E}[\rho(X,Y)]\leqslant D.$ (2)
Here, $I(X;Y)$ denotes the mutual information between the source $X$ and its reconstruction $Y$. To place rate-distortion theory on a rigorous mathematical footing, Csiszár established in his seminal work [11] a collection of fundamental properties of $R(D)$ and its parametric representations under remarkably weak assumptions. These results provide a foundational framework for source coding problems on general abstract spaces. We begin by imposing the most basic regularity condition on the distortion function itself.
Assumption 1.
Let $\rho:\mathcal{X}\times\mathcal{Y}\to[0,\infty]$ be a measurable function, referred to as the loss function. Assume that
$\inf_{y\in\mathcal{Y}}\rho(x,y)=0,\quad\forall x\in\mathcal{X}.$ (3)
This assumption ensures, in an intuitive sense, that for every source symbol $x$ there exists a sequence of reconstruction points whose distortion can be made arbitrarily small; in particular, there is no intrinsic positive distortion floor. Under this minimal condition, Csiszár first characterized the geometric properties of the rate-distortion function.
Lemma 1.
[Lemma 1.1 in [11]] Under Assumption 1, the rate-distortion function $R(D)$ is a non-increasing and convex function of $D$. Moreover, it admits the representation
$R(D)=\inf_{\mathbb{E}_{\pi}(\rho)=D}I(X;Y),$ (4)
for all $D\leqslant D_{1}$, where $D_{1}$ denotes the smallest nonnegative number such that $R(D)$ is constant on $(D_{1},+\infty)$ (typically corresponding to the threshold beyond which $R(D)=0$).
While the above lemma characterizes the shape of $R(D)$, it does not by itself guarantee that $R(D)$ is finite at any distortion level. To rule out the pathological situation in which $R(D)$ is identically infinite, an additional existence assumption is required.
Assumption 2.
Assume that there exists a reconstruction random variable $Y\sim\nu$ such that
$I(X;Y)<\infty,\quad\text{and}\quad\mathbb{E}\rho(X,Y)<\infty,$ (5)
where $X\sim\mu$ is the given source distribution.
This assumption guarantees the existence of at least one admissible coding scheme that achieves finite average distortion at finite mutual information cost. To accommodate more general topological structures, Csiszár further introduced a finite-approximation condition on the reproduction space.
Assumption 3.
Assume that there exists a finite set $B\subset\mathcal{Y}$ such that
$\int\rho(x,B)\,\mu(\mathrm{d}x)<\infty,\quad\text{where}\quad\rho(x,B)\coloneqq\min_{y\in B}\rho(x,y).$
Define
$D_{\min}\coloneqq\inf\{D\geqslant 0\mid R(D)<\infty\}$
as the smallest distortion level for which the rate-distortion function is finite. The following theorem provides sufficient conditions under which $R(D)$ is finite for all positive distortion levels, that is, $D_{\min}=0$.
Theorem 1.
[Theorem 2.1 in [11]] Under Assumptions 1 and 2, suppose that for every $\epsilon>0$ there exist measurable subsets $A_{1}\subset A_{2}\subset\cdots$ of $\mathcal{X}$ and finite subsets $B_{1}\subset B_{2}\subset\cdots$ of $\mathcal{Y}$ such that $\mu(A_{n})\to 1$ as $n\to\infty$, and
$\min_{y\in B_{n}}\rho(x,y)<\epsilon,\quad\forall x\in A_{n},\;n=1,2,\ldots$ (6)
Then $R(D)$ is finite for all $D>0$, i.e., $D_{\min}=0$.
The above condition is automatically satisfied when the reproduction alphabet $\mathcal{Y}$ is at most countable. More generally, it also holds if $\mathcal{X}$ is a Polish space and the truncated distortion function
$\rho^{(K)}(x,y)\coloneqq\min\{\rho(x,y),K\}$
is upper semi-continuous for some fixed $K>0$. In this case, since every probability measure on a Polish space is tight, one may choose the sets $A_{n}$ as an increasing sequence of compact subsets of $\mathcal{X}$ [11].
Having established the finiteness of $R(D)$, we next examine its asymptotic behavior at the extremes of the distortion range. The following theorem describes the convergence of $R(D)$ as the distortion constraint becomes arbitrarily large.
Theorem 2.
[Theorem 2.1 in [11]] Under Assumptions 1 and 2, suppose that there exist sets $A_{n}$ as in Theorem 1 and points $y_{n}\in\mathcal{Y}$ such that
$\max_{x\in A_{n}}\rho(x,y_{n})<\infty,\quad n=1,2,\ldots$ (7)
Then $R(D)\to 0$ as $D\to\infty$. Moreover, define the maximal distortion level
$D_{\max}=\lim_{K\to\infty}\left(\inf_{y}\int\rho^{(K)}(x,y)\,\mathrm{d}\mu\right).$ (8)
Then $R(D)>0$ for $D<D_{\max}$, and if $D_{\max}<\infty$, one has $R(D)=0$ for all $D>D_{\max}$.
Beyond qualitative existence and asymptotic properties, an explicit characterization of $R(D)$ is of central importance. By invoking Lagrange multiplier techniques, Csiszár derived a parametric representation of the rate-distortion function, revealing its dual relationship with convex optimization.
Theorem 3.
[Theorem 2.3 in [11]] Under Assumptions 1, 2, and 3, for every $D>D_{\min}$,
$R(D)=\max_{\beta,\alpha(x)}\left\{\int\log\alpha(x)\,\mathrm{d}\mu-\beta D\right\},$ (9)
where the maximization is taken over $\alpha(x)\geqslant 1$, $\beta\geqslant 0$, subject to the constraints
$\int\alpha(x)e^{-\beta\rho(x,y)}\,\mathrm{d}\mu\leqslant 1,\quad\forall y\in\mathcal{Y}.$ (10)
The maximizing value of $\beta$ is precisely the one associated with $D$ through the supporting hyperplane condition
$R(D^{\prime})+\beta D^{\prime}\geqslant R(D)+\beta D,\quad\forall D^{\prime}.$ (11)
The above results establish the theoretical backbone of the rate-distortion function. Building upon this framework, Gray provided a systematic reformulation and extension of Csiszár’s theory, recasting $R(D)$ into a compact variational form that directly underlies the Blahut-Arimoto algorithm and its generalizations.
Theorem 4.
[Theorem 9.2 in [18]] If $R(D)<\infty$, then the rate-distortion function admits the representation
$R(D)$ $=\max_{\beta\geqslant 0}\left\{F(\beta)-\beta D\right\},$ (12)
$F(\beta)$ $=\inf_{\nu\in\mathcal{P}(\mathcal{Y})}\int\log\frac{1}{\int e^{-\beta\rho(x,y)}\,\mathrm{d}\nu}\,\mathrm{d}\mu,$
where the infimum is taken over all probability measures $\nu$ on the reproduction space $\mathcal{Y}$. This formulation reduces the computation of $R(D)$ to a coupled optimization over the output distribution $\nu$ and the Lagrange multiplier $\beta$.
While the above variational representations provide deep structural insight into the rate-distortion function, they do not by themselves guarantee that the infima involved are attained by an optimal reconstruction distribution. Establishing such attainment—-especially in non-compact settings—-is essential for turning these representations into concrete optimality statements and for justifying the existence of minimizers. In classical treatments, such existence results are typically obtained under compactness assumptions on the reproduction space. A significant advance toward relaxing this restriction was achieved by Rezaei et al. [31], who developed a general existence theory for minimizers on non-compact spaces.
Theorem 5.
[Theorem 3.1 in [31]] Let $\mathcal{X}$ and $\mathcal{Y}$ be Polish spaces, and let $\rho:\mathcal{X}\times\mathcal{Y}\to[0,\infty]$ be a nonnegative extended-real-valued measurable distortion function. Let $\mu$ be a probability measure on the source space $\mathcal{X}$. Assume that for $\mu$-almost every $x\in\mathcal{X}$, the mapping $y\mapsto\rho(x,y)$ is continuous on $\mathcal{Y}$. If, for every $D\in[0,\infty)$, the admissible set is nonempty, then the rate-distortion problem admits a minimizer and an optimal reconstruction distribution exists.
Although the result of Rezaei et al. represents a substantial step forward by removing compactness assumptions on the reproduction space, Theorem 5 still relies on an additional regularity requirement, namely the pointwise continuity of the distortion function $y\mapsto\rho(x,y)$ for $\mu$-almost every $x$. As we demonstrate in this work, this continuity assumption is not intrinsic to the existence of optimal reconstructions and can be relaxed to the weaker requirement of lower semi-continuity.
In addition, to control the behavior of minimizing sequences in non-compact reproduction spaces, we impose a coercivity condition on the distortion function, which serves to prevent probability mass from escaping to infinity. Within this framework, it suffices to assume that the distortion function is proper, measurable, lower semi-continuous, and coercive. This setting generalizes the continuous case and extends Rezaei’s continuity-based existence theory.
In particular, our framework accommodates the following representative class of generalized distortion functions, which are discontinuous and hence cannot be handled by existing approaches directly:
  • Perceptual coding with JND thresholds (Dead-Zone): In perceptual image and video compression, the concept of Just Noticeable Difference (JND) implies that reconstruction errors below a certain visibility threshold $\tau$ are imperceptible to the human eye. An ideal perceptual distortion measure reflects this by imposing zero penalty within the "dead zone" and a standard penalty outside, which is also a hard thresholding scheme [13]. For instance, consider the distortion function:
    $\rho(x,y)=\|x-y\|^{2}\cdot\bm{1}_{\{\|x-y\|>\tau\}}.$
    This function exhibits a discontinuity at the threshold boundary $\|x-y\|=\tau$.
  • Robust distortion with penalty jumps: In anomaly detection and robust estimation, it is common to impose a large penalty once the reconstruction error exceeds a tolerance level. Furthermore, such functions also emerge in the field of discontinuous optimization [9, 27, 19, 16]. For example, consider the distortion function:
    $\rho(x,y)=\begin{cases}\left\|{x-y}\right\|^{2},&\text{if }\left\|{x-y}\right\|\leq\tau_{1},\\ \left\|{x-y}\right\|^{2}+J,&\text{if }\tau_{1}<\left\|{x-y}\right\|\leq\tau_{2},\\ +\infty,&\text{if }\left\|{x-y}\right\|>\tau_{2},\end{cases}$
    where $J$ is a large constant. The presence of abrupt penalty jumps renders such distortions incompatible with continuity assumptions.
By leveraging the concentration-compactness principle, we show that the existence of optimal reconstruction distributions can nevertheless be guaranteed under these broad conditions. This approach allows us to systematically rule out pathological minimizing sequences and to establish existence results in non-compact settings.
III Main Results
This section lays the analytical groundwork for our main existence theorems. Central to our approach is the concentration-compactness principle, a pivotal tool in nonlinear functional analysis for addressing variational problems where standard compactness arguments are insufficient. We introduce a generalized formulation, specifically adapted to the abstract Polish space setting pertinent to rate-distortion theory.
For the abstract rate-distortion problem considered in this paper, we require a generalized form of the above lemma that applies to probability measures on non-compact Polish spaces.
Theorem 6 (Generalized Concentration-compactness Lemma I).
Let $\{\mu_{m}\}$ be a sequence of probability measures on an unbounded Polish space $(\mathcal{X},d)$. Then there exists a subsequence (again denoted by $\{\mu_{m}\}$) such that one of the following alternatives holds:
  1. 1.
    Compactness: There exists a sequence $\{x_{m}\}\subset\mathcal{X}$ such that for every $\epsilon>0$, there exists $R>0$ with
    $\int_{B_{R}(x_{m})}\mathrm{d}\mu_{m}\geq 1-\epsilon,\quad\forall m\in\mathbb{N}.$
  2. 2.
    Vanishing: For every fixed $R>0$,
    $\lim_{m\to\infty}\left(\sup_{x\in\mathcal{X}}\int_{B_{R}(x)}\mathrm{d}\mu_{m}\right)=0.$
  3. 3.
    Dichotomy: There exists $\lambda\in(0,1)$ such that for every $\epsilon>0$, there exist $R>0$ and a sequence $\{x_{m}\}\subset\mathcal{X}$ with the property that for any $R^{\prime}>R$, there exist nonnegative measures $\mu_{m}^{1},\mu_{m}^{2}$ satisfying
    $0\leq\mu_{m}^{1}+\mu_{m}^{2}\leq\mu_{m},$
    $\mathrm{supp}(\mu_{m}^{1})\subset B_{R}(x_{m}),\quad\mathrm{supp}(\mu_{m}^{2})\subset B_{R^{\prime}}^{c}(x_{m}),$
    $\limsup_{m\to\infty}\Big(\big|\lambda-\mu_{m}^{1}(\mathcal{X})\big|+\big|(1-\lambda)-\mu_{m}^{2}(\mathcal{X})\big|\Big)\leq\epsilon,$
    $\mathrm{dist}\big(\mathrm{supp}(\mu_{m}^{1}),\mathrm{supp}(\mu_{m}^{2})\big)\to\infty.$ (13)
The proof is deferred to Appendix A, which follows the same lines as the classical version on $\mathbb{R}^{n}$.
Remark 1.
Here we emphasize that the Polish space $(\mathcal{X},d)$ in Theorem 6 is required to be unbounded, since for any bounded Polish space, the space itself is compact. In such cases, the tightness of probability measures can be established more directly via Prokhorov’s theorem, as demonstrated by Csiszár [11].
Remark 2.
We clarify the use of the term "compactness" within the context of Theorem 6. For general unbounded Polish spaces, this terminology represents a slight abuse of notation, as a closed ball $B_{R}(x_{m})$ is not necessarily a compact set in infinite-dimensional spaces. In our framework, "compactness" should be interpreted as the concentration of measure within a bounded set. Crucially, for this concentration to imply the tightness of the measure sequence, the underlying space must be proper.
Definition 1 (Proper Metric Spaces).
A metric space is called proper if every closed and bounded subset is compact.
This property ensures that a sequence of measures concentrated on a sequence of bounded sets is indeed tight, allowing for the application of Prokhorov’s theorem. For clarity, we provide several canonical examples of proper Polish spaces:
  • Any closed subset of $\mathbb{R}^{n}$ endowed with the Euclidean metric.
  • Any compact Polish space, such as $0,1]^{\mathbb{N}}$. This space is infinite-dimensional yet proper, as its compactness is guaranteed by Tychonoff’s theorem.
  • Every complete sub-Riemannian / sub-Finsler manifold is proper Polish, which is guaranteed by the Chow-Rashevskii theorem [8, 30] and the Ball-Box theorem [28, 22] assuming second-countability. For complete Riemannian / Finsler manifolds, this is a direct consequence of the Hopf-Rinow theorem [12].
Using the generalized concentration-compactness lemma, we can show that any minimizing sequence of the rate-distortion functional admits a tight subsequence. This property is the key analytical ingredient for establishing the existence of optimal reconstruction distributions.
Invoking the parametric representation of the rate-distortion function (see Theorem 4), we may rewrite $R(D)$ as
$R(D)$ $=$ $\inf_{\nu\in\mathcal{P}(\mathcal{Y})}\Bigg\{-\int_{\mathcal{X}}\log\!\left(\int e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu\right)\mathrm{d}\mu-\beta^{\star}D\Bigg\}$
$=$ $-\sup_{\nu\in\mathcal{P}(\mathcal{Y})}\Bigg\{\int_{\mathcal{X}}\log\!\left(\int e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu\right)\mathrm{d}\mu\Bigg\}+\beta^{\star}D,$
where $\beta^{\star}$ denotes the Lagrange multiplier associated with the distortion level $D$.
To ensure well-posedness of the above variational problem on non-compact spaces, we impose the following conditions on the distortion function, which is standard and essentially necessary in non-compact variational analysis.
Assumption 4 (Coercivity of the Distortion Function).
Let $\rho:\mathcal{X}\times\mathcal{Y}\to[0,\infty]$ be a lower semi-continuous proper measurable function that is coercive with respect to $y$, namely:
  1. 1.
    For $\mu$-almost every $x\in\mathcal{X}$, the mapping $y\mapsto\rho(x,y)$ is lower semi-continuous on $\mathcal{Y}$.
  2. 2.
    For each $x\in\mathcal{X}$, there exists $z\in\mathcal{Y}$ such that $\rho(x,z)<\infty$.
  3. 3.
    For any fixed $y^{\prime}\in\mathcal{Y}$,
    $\lim_{d_{\mathcal{Y}}(y^{\prime},y)\to\infty}\rho(x,y)=+\infty,\quad\forall x\in\mathcal{X}.$
Here $(\mathcal{Y},d_{\mathcal{Y}})$ is a Polish space, while $(\mathcal{X},d_{\mathcal{X}})$ is an arbitrary Polish space.
Remark 3 (On the Usage of the Term "Proper").
For the sake of clarity, we distinguish between the two standard mathematical meanings of the term proper used in Assumption 4 and Definition 1.
  • A function $f:\mathcal{X}\to[0,+\infty]$ is called proper if it is not identically equal to $\infty$.
  • When referring to $(\mathcal{Y},d_{\mathcal{Y}})$ as a proper Polish space, we use the term in its topological sense, meaning every closed and bounded subset is compact. This property was detailed in the preceding Remark 2 and is crucial for establishing the tightness of measure sequences.
These two conditions are independent and standard in their respective fields.
The lower semi-continuity ensures that the infimum in the variational problem is attainable when we take limits of minimizing sequences. This coercivity condition ensures that the distortion grows sufficiently fast as $y$ escapes to infinity, thereby preventing probability mass from dispersing or vanishing at infinity along minimizing sequences. From a physical perspective, it enforces an energetic penalty against reconstructions that drift arbitrarily far from the source. With these preparations, we are now ready to state our main existence theorem.
Theorem 7 (Tightness of Minimizing Sequences).
Let $\{\nu_{m}\}$ be a minimizing sequence for the rate-distortion functional $R(D)$, i.e., $\nu_{m}\in\mathcal{P}(\mathcal{Y})$ and
$R^{(m)}(D)\coloneqq-\int_{\mathcal{X}}\log\!\left(\int e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}\right)\mathrm{d}\mu-\beta^{\star}D.$
Assume that $\mathcal{Y}$ is a proper Polish space and that Assumption 4 holds. Then the sequence $\{\nu_{m}\}$ admits a tight subsequence, and consequently an optimal reconstruction distribution exists.
Sketch of Proof.
Without loss of generality, we assume that the space $\mathcal{Y}$ is unbounded; otherwise, the tightness of any sequence of probability measures is automatic by Prokhorov’s theorem, and the coercivity condition is not needed (see Remark 4). The detailed proof is provided in Appendix B. The argument proceeds in three main steps. First, we apply the Concentration-compactness Principle (Theorem 6) to a minimizing sequence $\{\nu_{m}\}$. Second, we show that the Vanishing and Dichotomy alternatives are incompatible with the minimizing property of the sequence, leading to a contradiction. Third, we deduce that the sequence must satisfy the Compactness alternative. By leveraging the coercivity of the distortion function, we show that this concentration of measure occurs within a bounded region of the space. The assumption that $\mathcal{Y}$ is a proper Polish space then allows us to conclude that the sequence is tight. The existence of a minimizer follows from Prokhorov’s theorem and the lower semi-continuity of the functional. ∎
Remark 4 (Relation to the Compact Case).
In the classical setting of a compact reproduction space $\mathcal{Y}$, the proof of existence simplifies considerably. The primary simplification arises from the fact that on a compact space, any sequence of probability measures is automatically tight by Prokhorov’s theorem. Consequently, the concentration-compactness principle and the coercivity condition (Assumption 4, item 3) become redundant.
However, another crucial step is to ensure that the weak limit of a minimizing subsequence is indeed a minimizer. This requires the lower semi-continuity of the rate-distortion functional, which can be established via lower semi-continuity of $\rho(x,y)$ with respect to $y$. This relaxes the stronger continuity assumptions that are often presumed in classical treatments.
IV Discussion and Future Work
In this paper, we established the existence of optimal reconstruction distributions for rate-distortion problems on general proper Polish spaces. By introducing the concentration-compactness principle and leveraging coercivity conditions, our framework provides a transparent variational approach that unifies and extends existing existence results.
The analytical framework presented here opens up several avenues for future research. A primary direction is to address models with additional structural constraints, which are common in applications. For instance, in non-anticipative (or causal) rate-distortion theory [5, 7, 36, 6], existence results on non-compact spaces remain scarce, largely due to the analytical challenges posed by causal constraints. Adapting a variational approach to rigorously establish the existence of optimal causal coders in such non-compact settings remains a significant open problem for future investigation.
Appendix A Proof of Theorem 6
The proof follows the classical strategy of [23, 24, 37], adapted to the abstract Polish-space setting. We begin by introducing the notion of the concentration function associated with a measure $\mu$:
$Q(r)\coloneqq\sup_{x\in\mathcal{X}}\int_{B_{r}(x)}\,\mathrm{d}\mu.$ (14)
Let $Q_{m}$ denote the concentration function associated with the measure $\mu_{m}$. Observe that $\{Q_{m}\}$ is a sequence of non-decreasing, nonnegative, and bounded functions on $\mathbb{R}_{+}$. Moreover, since each $\mu_{m}$ is a probability measure, it holds that
$\lim_{R\to\infty}Q_{m}(R)=1.$
Complementing this principle, our analysis will also leverage a classical result from real analysis, which we state below for completeness.
Lemma 2.
[Helly’s Selection Theorem] Let $\{f_{n}\}$ be a sequence of nondecreasing real-valued functions on an interval $I\subset\mathbb{R}$. Assume that the sequence is uniformly bounded, i.e., there exist constants $a,b\in\mathbb{R}$ such that $a\leq f_{n}\leq b,\forall n$. Then there exists a subsequence that converges pointwise on $I$.
By Helly’s selection theorem, there exists a subsequence $\{Q_{m}\}$ (by a slight abuse of notation) that converges pointwise for almost every $R>0$ to a non-decreasing, nonnegative, bounded function $Q$. Since any monotone function admits left and right limits and has at most countably many discontinuities, we normalize $Q$ to be left-continuous by defining
$Q(R)=\lim_{\epsilon\to 0^{+}}Q(R-\epsilon).$
Fix an arbitrary $R>0$. We consider a sequence $\{R_{k}\}$ such that $R_{k}\nearrow R$ and
$\lim_{m\to\infty}Q_{m}(R_{k})=Q(R_{k}).$
Since each $Q_{m}$ is non-decreasing, we have $Q_{m}(R_{k})\leq Q_{m}(R)$ for all $k,m\in\mathbb{N}$. Consequently,
$\liminf_{m\to\infty}Q_{m}(R_{k})=\lim_{m\to\infty}Q_{m}(R_{k})=Q(R_{k})\leq\liminf_{m\to\infty}Q_{m}(R).$
Letting $R_{k}\to R$ and using the left-continuity of $Q$, we obtain
$\lim_{k\to\infty}Q(R_{k})=Q(R)\leq\liminf_{m\to\infty}Q_{m}(R).$
Define
$\lambda\coloneqq\lim_{R\to\infty}Q(R),$
which clearly satisfies $\lambda\in[0,1]$. We now distinguish three cases.
Case 1: $\lambda=0$ (Vanishing). In this case, by definition of $\lambda$, we have
$\lim_{m\to\infty}\left(\sup_{x\in\mathcal{X}}\int_{B_{R}(x)}\,\mathrm{d}\mu_{m}\right)=0,\qquad\forall R>0,$
which is precisely the vanishing alternative.
Case 2: $\lambda=1$ (Compactness). Then there exists some $R_{0}>0$ such that $Q(R_{0})>\tfrac{1}{2}$. For each $m\in\mathbb{N}$, choose $x_{m}\in\mathcal{X}$ satisfying
$Q_{m}(R_{0})\leq\int_{B_{R_{0}}(x_{m})}\,\mathrm{d}\mu_{m}+\frac{1}{m},\qquad\lim_{m\to\infty}Q_{m}(R_{0})=Q(R_{0}).$
Fix $0<\epsilon<\tfrac{1}{2}$. Choose $R>0$ such that $Q(R)>1-\epsilon>\tfrac{1}{2}$. For each $m$, let $y_{m}\in\mathcal{X}$ satisfy
$Q_{m}(R)\leq\int_{B_{R}(y_{m})}\,\mathrm{d}\mu_{m}+\frac{1}{m},\qquad\lim_{m\to\infty}Q_{m}(R)=Q(R),$
where the choice of $y_{m}$ may depend on $\epsilon$. Then
$\int_{B_{R_{0}}(x_{m})}\,\mathrm{d}\mu_{m}+\int_{B_{R}(y_{m})}\,\mathrm{d}\mu_{m}$ $\geq$ $Q_{m}(R_{0})+Q_{m}(R)-\frac{2}{m}$ (15)
$\to$ $Q(R_{0})+Q(R)$
$>$ $\frac{1}{2}+\frac{1}{2}=1=\int_{\mathcal{X}}\,\mathrm{d}\mu_{m},$
for all sufficiently large $m$. Hence the balls $B_{R_{0}}(x_{m})$ and $B_{R}(y_{m})$ must intersect. Otherwise,
$\int_{B_{R_{0}}(x_{m})}\,\mathrm{d}\mu_{m}+\int_{B_{R}(y_{m})}\,\mathrm{d}\mu_{m}$ $=$ $\int_{B_{R_{0}}(x_{m})\cup B_{R}(y_{m})}\,\mathrm{d}\mu_{m}$ (16)
$\leq$ $\int_{\mathcal{X}}\,\mathrm{d}\mu_{m}=1,$
which leads to a contradiction.
Therefore,
$\mathrm{dist}(x_{m},y_{m})\leq R_{0}+R,\qquad B_{R}(y_{m})\subset B_{2R+R_{0}}(x_{m}),$
$\int_{B_{2R+R_{0}}(x_{m})}\,\mathrm{d}\mu_{m}\geq\int_{B_{R}(y_{m})}\,\mathrm{d}\mu_{m}\geq Q_{m}(R)-\frac{1}{m}\geq 1-\epsilon,$ (17)
for all sufficiently large $m$. Thus there exists $M\in\mathbb{N}$ such that the above holds for all $m\geq M$. Moreover, since finitely many indices are involved, there exists $\widetilde{R}>0$ such that
$\int_{B_{\widetilde{R}}(x_{m})}\,\mathrm{d}\mu_{m}\geq 1-\epsilon,\qquad\forall\,1\leq m\leq M.$
Setting $R^{\prime}=\max\{\widetilde{R},\,2R+R_{0}\}$ yields the compactness alternative.
Case 3: $0<\lambda<1$ (Dichotomy). Fix $\epsilon>0$. There exists $R(\epsilon)>0$ such that
$Q(R)>\lambda-\frac{\epsilon}{3},\qquad\lim_{m\to\infty}Q_{m}(R)=Q(R).$
Hence there exists $M(\epsilon)\in\mathbb{N}$ such that
$Q_{m}(R)>Q(R)-\frac{\epsilon}{3}>\lambda-\frac{2\epsilon}{3},\qquad\forall m\geq M(\epsilon).$
Choose a sequence $\{x_{m}\}$ depending on $\epsilon$ such that
$Q_{m}(R)\geq\int_{B_{R}(x_{m})}\,\mathrm{d}\mu_{m}\geq Q_{m}(R)-\frac{\epsilon}{3}>\lambda-\epsilon.$
Similarly, there exist $R^{\prime}(\epsilon)>0$ and $M^{\prime}(\epsilon)\geq M(\epsilon)$ such that
$\lambda-\epsilon<Q_{m}(R^{\prime})<\lambda+\epsilon,\qquad\forall m\geq M^{\prime}(\epsilon).$
Since $\lim_{r\to\infty}Q(r)=\lambda$, there exists $R_{0}<\infty$ such that
$Q(r)<\lambda+\frac{\epsilon}{2},\qquad\forall r\geq R_{0}.$
Choose a strictly increasing sequence $\{R_{k}\}_{k\geq 1}$ with $R_{k}\to\infty$ and $R_{1}\geq R_{0}$, such that
$Q(R_{k})<\lambda+\frac{\epsilon}{2},\qquad\lim_{m\to\infty}Q_{m}(R_{k})=Q(R_{k}),\quad\forall k\in\mathbb{N}.$
By a diagonal construction, we define an increasing index sequence $\{M_{k}\}_{k\geq 1}$ such that
$Q_{m}(R_{k})<Q(R_{k})+\frac{\epsilon}{2}<\lambda+\epsilon,\qquad\forall m\geq M_{k}.$
Using $\{R_{k}\}$ and $\{M_{k}\}$, define a sequence $\{\widetilde{R}_{m}\}$ by
$\widetilde{R}_{m}=\begin{cases}R_{0},&M^{\prime}\leq m<M_{1},\\ R_{k},&M_{k}\leq m<M_{k+1}.\end{cases}$
In summary, for any $\epsilon>0$, there exist $R>0$, $m_{0}\in\mathbb{N}$, a sequence $\{R_{m}\}$ with $R_{m}\to\infty$, and points $\{x_{m}\}\subset\mathcal{X}$ such that for all $m\geq m_{0}$,
$Q_{m}(R)\geq\int_{B_{R}(x_{m})}\,\mathrm{d}\mu_{m}>\lambda-\epsilon,$ (18)
$Q_{m}(R)\leq Q_{m}(R_{m})<\lambda+\epsilon.$ (19)
For any $R^{\prime}>R$, we may increase $m_{0}$ so that $R_{m}\geq R^{\prime}$ for all $m\geq m_{0}$. Define
$\mu_{m}^{1}=\mu_{m}\cdot\bm{1}_{B_{R}(x_{m})},\qquad\mu_{m}^{2}=\mu_{m}\cdot\bm{1}_{\mathcal{X}\setminus B_{R_{m}}(x_{m})}.$
Then $0\leq\mu_{m}^{1}+\mu_{m}^{2}\leq\mu_{m}$, with
$\mathrm{supp}(\mu_{m}^{1})\subset B_{R}(x_{m}),\qquad\mathrm{supp}(\mu_{m}^{2})\subset\mathcal{X}\setminus B_{R_{m}}(x_{m})\subset\mathcal{X}\setminus B_{R^{\prime}}(x_{m}).$
Finally, for all $m\geq m_{0}$,
$\left|\lambda-\int_{\mathcal{X}}\,\mathrm{d}\mu_{m}^{1}\right|+\left|(1-\lambda)-\int_{\mathcal{X}}\,\mathrm{d}\mu_{m}^{2}\right|=\left|\lambda-\int_{B_{R}(x_{m})}\,\mathrm{d}\mu_{m}\right|+\left|\lambda-\int_{B_{R_{m}}(x_{m})}\,\mathrm{d}\mu_{m}\right|<2\epsilon.$ (20)
This establishes the dichotomy alternative and completes the proof.
Appendix B Proof of Theorem 7
As a preliminary step, we first establish the lower semi-continuity of the rate-distortion functional with respect to the weak topology on the space of probability measures.
Lemma 3 (Lower Semi-continuity of the Rate-distortion Functional).
Let $\mathcal{X}$ and $\mathcal{Y}$ be Polish spaces, and let $\mu\in\mathcal{P}(\mathcal{X})$ be a fixed source distribution. Suppose the distortion function $\rho:\mathcal{X}\times\mathcal{Y}\to[0,\infty]$ satisfies the following conditions:
  1. 1.
    For every $y\in\mathcal{Y}$, the map $x\mapsto\rho(x,y)$ is measurable.
  2. 2.
    For $\mu$-almost every $x\in\mathcal{X}$, the map $y\mapsto\rho(x,y)$ is lower semi-continuous on $\mathcal{Y}$.
Then, for any $\beta\geq 0$, the dual functional $J_{\beta}:\mathcal{P}(\mathcal{Y})\to\mathbb{R}\cup\{+\infty\}$ defined by
$J_{\beta}(\nu)\coloneqq-\int_{\mathcal{X}}\log\left(\int_{\mathcal{Y}}e^{-\beta\rho(x,y)}\,\mathrm{d}\nu(y)\right)\mathrm{d}\mu(x)$ (21)
is lower semi-continuous with respect to the weak topology on $\mathcal{P}(\mathcal{Y})$. That is, if $\nu_{m}\xrightarrow{w}\nu$, then $J_{\beta}(\nu)\leq\liminf\limits_{m\to\infty}J_{\beta}(\nu_{m})$.
Proof.
Let $\{\nu_{m}\}_{m=1}^{\infty}$ be a sequence in $\mathcal{P}(\mathcal{Y})$ converging weakly to $\nu$. We proceed in three steps.
Step 1: Upper semi-continuity of the inner integral. Fix $x\in\mathcal{X}$ and $\beta\geq 0$. Define the function $h_{x}:\mathcal{Y}\to[0,1]$ by $h_{x}(y)\coloneqq e^{-\beta\rho(x,y)}$. Since $\rho(x,\cdot)$ is lower semi-continuous, the function $h_{x}(\cdot)$ is upper semi-continuous and bounded. According to the Portmanteau theorem, weak convergence $\nu_{m}\xrightarrow{w}\nu$ implies
$\limsup_{m\to\infty}\int_{\mathcal{Y}}h_{x}(y)\,\mathrm{d}\nu_{m}(y)\leq\int_{\mathcal{Y}}h_{x}(y)\,\mathrm{d}\nu(y).$ (22)
Step 2: Lower semi-continuity of the integrand. Let $\Phi_{m}(x)\coloneqq\int_{\mathcal{Y}}e^{-\beta\rho(x,y)}\,\mathrm{d}\nu_{m}(y)$ and $\Phi(x)\coloneqq\int_{\mathcal{Y}}e^{-\beta\rho(x,y)}\,\mathrm{d}\nu(y)$. The inequality from Step 1 can be rewritten as $\limsup\limits_{m\to\infty}\Phi_{m}(x)\leq\Phi(x)$. Since the function $t\mapsto-\log(t)$ is continuous and monotonically decreasing on $(0,1]$, we have
$\liminf_{m\to\infty}\big(-\log\Phi_{m}(x)\big)=-\log\left(\limsup_{m\to\infty}\Phi_{m}(x)\right)\geq-\log\Phi(x).$ (23)
This inequality holds even if $\Phi(x)=0$, in which case both sides are $\infty$. Thus, the integrand in $J_{\beta}$ is pointwise lower semi-continuous with respect to $\nu$.
Step 3: Application of Fatou’s Lemma. Since the integrand $\log\Phi_{m}(x)$ is bounded from below (specifically, $\log\Phi_{m}(x)\geq 0$ if we assume normalized measures or simply note that the integral is bounded by 1), we can apply Fatou’s Lemma to the outer integral with respect to $\mu$:
$\liminf_{m\to\infty}J_{\beta}(\nu_{m})$ $=\liminf_{m\to\infty}\int_{\mathcal{X}}-\log\Phi_{m}(x)\,\mathrm{d}\mu(x)$ (24)
$\geq\int_{\mathcal{X}}\liminf_{m\to\infty}\big(-\log\Phi_{m}(x)\big)\,\mathrm{d}\mu(x)$
$\geq\int_{\mathcal{X}}-\log\Phi(x)\,\mathrm{d}\mu(x)$
$=J_{\beta}(\nu).$
This confirms that $J_{\beta}(\nu)$ is lower semi-continuous. ∎
Now we proceed to prove Theorem 7. Without loss of generality, we assume that $\beta^{\star}>0$ and that $\{\nu_{m}\}$ is a monotonically decreasing minimizing sequence, i.e., $R^{(m+1)}(D)\leq R^{(m)}(D)$ for every $m\in\mathbb{N}$. Additionally, we assume that the space $\mathcal{Y}$ is unbounded; otherwise, the tightness of any sequence of probability measures is automatic by Prokhorov’s theorem, and the coercivity condition is not needed (see Remark 4).
Consequently, we have $\lim\limits_{m\to\infty}R^{(m)}(D)=R(D)$. By Theorem 6, there exists a subsequence $\{\nu_{m}\}$ satisfying one of the following three alternatives:
  1. 1.
    Compactness: There exists a sequence $\{y_{m}\}\subset\mathcal{Y}$ such that for every $\epsilon>0$, there exists $R>0$ with
    $\int_{B_{R}(y_{m})}\mathrm{d}\nu_{m}\geq 1-\epsilon,\quad\forall m\in\mathbb{N}.$
  2. 2.
    Vanishing: For every fixed $R>0$, it holds that
    $\lim_{m\to\infty}\left(\sup_{y\in\mathcal{Y}}\int_{B_{R}(y)}\mathrm{d}\nu_{m}\right)=0.$
  3. 3.
    Dichotomy: There exists $\lambda\in(0,1)$ such that for every $\epsilon>0$, there exist $R>0$ and a sequence $\{y_{m}\}\subset\mathcal{Y}$ with the following property: Given any $R^{\prime}>R$, there exist non-negative measures $\nu_{m}^{1},\nu_{m}^{2}$ such that
    $0\leqslant\nu_{m}^{1}+\nu_{m}^{2}\leqslant\nu_{m},$ (25)
    $\mathrm{supp}(\nu_{m}^{1})\subset B_{R}(y_{m}),\quad\mathrm{supp}(\nu_{m}^{2})\subset B_{R^{\prime}}^{c}(y_{m}),$
    $\limsup_{m\to\infty}\left(\left|\lambda-\int_{\mathcal{Y}}\mathrm{d}\nu_{m}^{1}\right|+\left|(1-\lambda)-\int_{\mathcal{Y}}\mathrm{d}\nu_{m}^{2}\right|\right)\leqslant\epsilon,$
    $\mathrm{dist}(\mathrm{supp}(\nu_{m}^{1}),\mathrm{supp}(\nu_{m}^{2}))\to\infty,\quad\text{as }m\to\infty.$
Our proof strategy is to demonstrate that any minimizing sequence for the rate-distortion functional must be tight. We achieve this by applying the Generalized Concentration-compactness Principle (Theorem 6) and showing that both the Vanishing and Dichotomy alternatives lead to contradictions, leaving Compactness as the only possibility. The tightness then follows from the Compactness alternative, combined with the coercivity of the distortion function and the properness of the reproduction space.
Step 1: Ruling out Vanishing.
First, we show that the Vanishing alternative is incompatible with the minimizing property of the sequence. Assume, for the sake of contradiction, that the sequence $\{\nu_{m}\}$ vanishes. This implies that for any fixed radius $R>0$,
$\lim_{m\to\infty}\left(\sup_{y\in\mathcal{Y}}\nu_{m}(B_{R}(y))\right)=0.$
Let us fix an $x\in\mathcal{X}$ for which the coercivity condition holds and a reference point $y_{0}\in\mathcal{Y}$. By the coercivity of $\rho$ (Assumption 4), for any arbitrarily large constant $M>0$, there exists a radius $K_{M}>0$ such that $\rho(x,y)\geq M$ for all $y\in\mathcal{Y}\setminus B_{K_{M}}(y_{0})$. We can then bound the inner integral of the functional $J_{\beta^{\star}}$ as follows:
$\int_{\mathcal{Y}}e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}(y)$ $=\int_{B_{K_{M}}(y_{0})}e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}(y)+\int_{\mathcal{Y}\setminus B_{K_{M}}(y_{0})}e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}(y)$
$\leq 1\cdot\nu_{m}(B_{K_{M}}(y_{0}))+e^{-\beta^{\star}M}\cdot\nu_{m}(\mathcal{Y}\setminus B_{K_{M}}(y_{0}))$
$\leq\nu_{m}(B_{K_{M}}(y_{0}))+e^{-\beta^{\star}M}.$
Taking the limit superior as $m\to\infty$, the first term vanishes due to the Vanishing property, yielding
$\limsup_{m\to\infty}\int_{\mathcal{Y}}e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}(y)\leq e^{-\beta^{\star}M}.$
Since this inequality holds for any arbitrarily large $M$, we can conclude that the limit must be zero. This pointwise convergence for $\mu$-almost every $x$ allows us to apply Fatou’s Lemma to the full functional:
$R(D)=\lim_{m\to\infty}J_{\beta^{\star}}(\nu_{m})$ $=\liminf_{m\to\infty}\int_{\mathcal{X}}\left(-\log\int_{\mathcal{Y}}e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}(y)\right)\mathrm{d}\mu(x)$
$\geq\int_{\mathcal{X}}\liminf_{m\to\infty}\left(-\log\int_{\mathcal{Y}}e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}(y)\right)\mathrm{d}\mu(x).$
As the inner integral tends to 0, its negative logarithm tends to $\infty$. Consequently, the outer integral diverges, implying $R(D)=+\infty$. This contradicts our initial assumption that $R(D)$ is finite. Therefore, the Vanishing case is ruled out.
Step 2: Ruling out Dichotomy.
Next, we preclude the Dichotomy alternative. We can construct a subsequence $\{\nu_{m}\}$ satisfying the following properties:
$0\leqslant\nu_{m}^{1}+\nu_{m}^{2}\leqslant\nu_{m},$ (26)
$\mathrm{supp}(\nu_{m}^{1})\subset B_{R}(y_{m_{k}}),\quad\mathrm{supp}(\nu_{m}^{2})\subset B_{R^{\prime}}^{c}(y_{m_{k}}),$
$\limsup_{k\to\infty}\left(\left|\lambda-\int_{\mathcal{Y}}\mathrm{d}\nu_{m}^{1}\right|+\left|(1-\lambda)-\int_{\mathcal{Y}}\mathrm{d}\nu_{m}^{2}\right|\right)\leqslant\epsilon,$
$\mathrm{dist}(\mathrm{supp}(\nu_{m}^{1}),\mathrm{supp}(\nu_{m}^{2}))\to\infty,\quad\text{as }k\to\infty.$
This allows for the decomposition:
$\int_{\mathcal{Y}}e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}$ $=\int_{\mathrm{supp}(\nu_{m}^{1})}e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}+\int_{\mathrm{supp}(\nu_{m}^{2})}e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}+\int_{\mathcal{Y}\backslash\left(\mathrm{supp}(\nu_{m}^{1})\cup\mathrm{supp}(\nu_{m}^{2})\right)}e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}$ (27)
$=\int_{\mathcal{Y}}e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}^{1}+\int_{\mathcal{Y}}e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}^{2}+\int_{\mathcal{Y}}e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}^{3},$
where $\nu_{m}^{3}=\nu_{m}-\nu_{m}^{1}-\nu_{m}^{2}$. We then construct a modified measure $\tilde{\nu}_{m}$ by delicately tuning the components of $\nu_{m}$:
$\tilde{\nu}_{m}\coloneqq\alpha_{m}\cdot\nu_{m}^{1}+\beta_{m}\cdot\nu_{m}^{2}+\nu_{m}^{3},$ (28)
where $\alpha_{m}\int\mathrm{d}\nu_{m}^{1}+\beta_{m}\int\mathrm{d}\nu_{m}^{2}+\int\mathrm{d}\nu_{m}^{3}=1$, with $\alpha_{m},\beta_{m}>0$. For sufficiently large $m$, the properties of Dichotomy imply:
$\alpha_{m}(\lambda-\epsilon)+\beta_{m}(1-\lambda-\epsilon)-\epsilon\leq 1\leq\alpha_{m}(\lambda+\epsilon)+\beta_{m}(1-\lambda+\epsilon)+\epsilon.$ (29)
Setting $\alpha_{m}\equiv\dfrac{1+\lambda^{-1}}{2}=\dfrac{1+\lambda}{2\lambda}$ for large $m$, we ensure that $\beta_{m}$ is bounded strictly away from zero and infinity:
$\dfrac{1-\epsilon-\frac{(\lambda+1)(\lambda+\epsilon)}{2\lambda}}{1-\lambda+\epsilon}\leq\beta_{m}\leq\dfrac{1+\epsilon-\frac{(\lambda+1)(\lambda-\epsilon)}{2\lambda}}{1-\lambda-\epsilon}.$ (30)
By choosing $\epsilon$ sufficiently small, $\beta_{m}$ can be constrained within a compact interval $\frac{1}{2}-\delta,\frac{1}{2}+\delta]$ for all large $m$, where $\delta\ll 1$.
For every large $m$ and fixed $x\in\mathcal{X}$, we estimate the difference between the integrals associated with $\nu_{m}$ and $\tilde{\nu}_{m}$ as follows:
$-\log\left(\int e^{-\beta^{\star}\rho(x,y)}d\tilde{\nu}_{m}\right)$ $-(-1)\cdot\log\left(\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}\right)$ (31)
$=\log\dfrac{\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}^{1}+\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}^{2}+\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}^{3}}{\alpha_{m}\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}^{1}+\beta_{m}\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}^{2}+\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}^{3}}$
$<\log\dfrac{\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}^{1}+2\epsilon}{\frac{1+\lambda}{2\lambda}\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}^{1}-(2^{-1}+\delta+1)\epsilon}$
$=-\log\left(1+\dfrac{1-\lambda}{2\lambda}\right)+\log\dfrac{\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}^{1}+2\epsilon}{\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}^{1}-\dfrac{(3+2\delta)\lambda}{1+\lambda}\epsilon}.$
Without loss of generality, we assume $\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}^{1}>0$, which is justified by the properness assumption on the loss function $\rho$. We can then choose $\epsilon$ sufficiently small such that the second term in the above inequality is bounded by a small constant $\epsilon^{\prime}$, yielding:
$-\log\left(\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\tilde{\nu}_{m}\right)$ $-(-1)\cdot\log\left(\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}\right)<-\log\left(1+\dfrac{1-\lambda}{2\lambda}\right)+\epsilon^{\prime}<-\log\left(1+\dfrac{1-\lambda}{3\lambda}\right).$ (32)
Applying Fatou’s Lemma, we obtain the following estimate:
$\liminf_{m\to\infty}\left(R^{(m)}(D)-\tilde{R}^{(m)}(D)\right)$ $=\liminf_{m\to\infty}\int_{\mathcal{X}}\left\{-\log\left(\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}\right)+\log\left(\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\tilde{\nu}_{m}\right)\right\}\mathrm{d}\mu$ (33)
$\geq\int_{\mathcal{X}}\liminf_{m\to\infty}\left\{-\log\left(\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\nu_{m}\right)+\log\left(\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\tilde{\nu}_{m}\right)\right\}\mathrm{d}\mu$
$>\log\left(1+\dfrac{1-\lambda}{3\lambda}\right)>0,$
where
$\tilde{R}^{(m)}(D)\coloneqq-\int_{\mathcal{X}}\log\left(\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\tilde{\nu}_{m}\right)\mathrm{d}\mu-\beta^{\star}D.$
Consequently, for sufficiently large $m$, there exists a measure $\tilde{\nu}_{m}$ such that
$\tilde{R}^{(m)}(D)$ $\coloneqq-\int_{\mathcal{X}}\log\left(\int e^{-\beta^{\star}\rho(x,y)}\mathrm{d}\tilde{\nu}_{m}\right)\mathrm{d}\mu-\beta^{\star}D$ (34)
$\leq R^{(m)}(D)-\log\left(1+\dfrac{1-\lambda}{3\lambda}\right)$
$\leq R(D)-\log\left(1+\dfrac{1-\lambda}{4\lambda}\right)$
$<R(D),$
which contradicts the assumption that $\{\nu_{m}\}$ is a minimizing sequence for $R(D)$. Thus, the Dichotomy case is also ruled out.
Step 3: Concluding with Compactness, Tightness, and Existence.
Since Vanishing and Dichotomy have been excluded, the minimizing sequence $\{\nu_{m}\}$ must satisfy the Compactness alternative. This means there exists a sequence of centers $\{y_{m}\}\subset\mathcal{Y}$ such that for any $\epsilon>0$, there is a radius $R>0$ for which
$\nu_{m}(B_{R}(y_{m}))\geq 1-\epsilon,\quad\forall m\in\mathbb{N}.$
This property establishes that the sequence of measures is concentrated on a sequence of uniformly bounded sets, but it does not, by itself, prevent these sets from escaping to infinity.
At this juncture, we leverage the key assumptions of the theorem to show that the sequence of centers $\{y_{m}\}$ must be bounded. We argue by contradiction. Assume that $\{y_{m}\}$ is unbounded. Then there exists a subsequence, which we relabel as $\{y_{m}\}$ for notational simplicity, such that $d(y_{m},y_{0})\to\infty$ as $m\to\infty$ for some fixed reference point $y_{0}\in\mathcal{Y}$.
Let us analyze the consequence of this unboundedness on the inner integral of the functional $J_{\beta^{\star}}$. By the coercivity of $\rho$ (Assumption 4), for any $x\in\mathcal{X}$ (for which the condition holds) and any arbitrarily large constant $M>0$, there exists a radius $K_{M}>0$ such that
$\rho(x,y)\geq M\quad\text{for all }y\in\mathcal{Y}\setminus B_{K_{M}}(y_{0}).$ (35)
Since $d(y_{m},y_{0})\to\infty$, for the fixed radius $R$ from the Compactness alternative, there must exist an integer $N_{M}$ such that for all $m\geq N_{M}$, we have $d(y_{m},y_{0})>K_{M}+R$. By the triangle inequality, this ensures that the ball $B_{R}(y_{m})$ is completely contained in the coercive region, i.e.,
$B_{R}(y_{m})\subset\mathcal{Y}\setminus B_{K_{M}}(y_{0})\quad\text{for all }m\geq N_{M}.$
This means that for any $y\in B_{R}(y_{m})$, the bound from Eq. (35) applies.
We estimate the inner integral for $m\geq N_{M}$ by splitting the domain of integration, analogous to the Step 1:
$\int_{\mathcal{Y}}e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}(y)$ $=\int_{B_{R}(y_{m})}e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}(y)+\int_{\mathcal{Y}\setminus B_{R}(y_{m})}e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}(y).$
For the first term, since $y\in B_{R}(y_{m})$ implies $\rho(x,y)\geq M$, we have $e^{-\beta^{\star}\rho(x,y)}\leq e^{-\beta^{\star}M}$. For the second term, we use the trivial bound $e^{-\beta^{\star}\rho(x,y)}\leq 1$. This gives:
$\int_{\mathcal{Y}}e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}(y)$ $\leq e^{-\beta^{\star}M}\cdot\nu_{m}(B_{R}(y_{m}))+1\cdot\nu_{m}(\mathcal{Y}\setminus B_{R}(y_{m})).$
From the Compactness alternative, we know that for any $\epsilon>0$, there exists an $R$ such that $\nu_{m}(B_{R}(y_{m}))\geq 1-\epsilon$ and consequently $\nu_{m}(\mathcal{Y}\setminus B_{R}(y_{m}))<\epsilon$. Substituting these into the inequality, we get:
$\int_{\mathcal{Y}}e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}(y)\leq(1-\epsilon)e^{-\beta^{\star}M}+\epsilon.$
This inequality holds for all $m\geq N_{M}$. Taking the limit superior as $m\to\infty$, we find
$\limsup_{m\to\infty}\int_{\mathcal{Y}}e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}(y)\leq(1-\epsilon)e^{-\beta^{\star}M}+\epsilon.$
Since this is true for any $\epsilon>0$, we can let $\epsilon\to 0$, which yields
$\limsup_{m\to\infty}\int_{\mathcal{Y}}e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}(y)\leq e^{-\beta^{\star}M}.$
Furthermore, since this holds for any arbitrarily large $M>0$, we can conclude that the limit must be zero:
$\lim_{m\to\infty}\int_{\mathcal{Y}}e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}(y)=0.$
This pointwise convergence to zero for the inner integral for $\mu$-almost every $x$ allows us to apply Fatou’s Lemma to the full functional:
$\lim_{m\to\infty}J_{\beta^{\star}}(\nu_{m})$ $=\lim_{m\to\infty}\int_{\mathcal{X}}\left(-\log\int_{\mathcal{Y}}e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}(y)\right)\mathrm{d}\mu(x)$
$\geq\int_{\mathcal{X}}\liminf_{m\to\infty}\left(-\log\int_{\mathcal{Y}}e^{-\beta^{\star}\rho(x,y)}\,\mathrm{d}\nu_{m}(y)\right)\mathrm{d}\mu(x).$
Since the argument of the logarithm tends to 0, the integrand tends to $\infty$. Thus,
$\lim_{m\to\infty}J_{\beta^{\star}}(\nu_{m})=+\infty.$
This contradicts our initial assumption that $\{\nu_{m}\}$ is a minimizing sequence, which must converge to the finite value $R(D)$. Therefore, our premise that $\{y_{m}\}$ is unbounded must be false. The sequence of centers $\{y_{m}\}$ is bounded.
Since $\{y_{m}\}$ is bounded, there exists a larger, fixed ball $B_{R^{\prime}}(y_{0})$ that contains all the balls $B_{R}(y_{m})$. This implies that for any $\epsilon>0$, there exists a fixed bounded set $B^{\prime}=B_{R^{\prime}}(y_{0})$ such that
$\nu_{m}(B^{\prime})\geq 1-\epsilon,\quad\forall m\in\mathbb{N}.$
Next, we invoke the assumption that $\mathcal{Y}$ is a proper Polish space. By definition of a proper space, any closed and bounded set is compact. Thus, we have found a fixed compact set $K=B^{\prime}$ such that $\nu_{m}(K)\geq 1-\epsilon$ for all $m$. This is precisely the definition of tightness for the sequence $\{\nu_{m}\}$. By Prokhorov’s Theorem, the tightness of $\{\nu_{m}\}$ implies that its closure in the space of probability measures $\mathcal{P}(\mathcal{Y})$ is compact with respect to the weak topology. Therefore, there exists a subsequence $\{\nu_{m_{k}}\}$ and a measure $\nu\in\mathcal{P}(\mathcal{Y})$ such that $\nu_{m_{k}}\xrightarrow{w}\nu$ as $k\to\infty$.
Finally, by the lower semi-continuity of the functional $J_{\beta^{\star}}$ (established in Lemma 3) and the fact that $\{\nu_{m_{k}}\}$ is a minimizing sequence, we have:
$J_{\beta^{\star}}(\nu)\leq\liminf_{k\to\infty}J_{\beta^{\star}}(\nu_{m_{k}})=\lim_{m\to\infty}J_{\beta^{\star}}(\nu_{m})=R(D).$
Since $R(D)$ is defined as the infimum over all measures, we must also have $J_{\beta^{\star}}(\nu)\geq R(D)$. Combining these inequalities, we get $J_{\beta^{\star}}(\nu)=R(D)$. This shows that the infimum is attained at $\nu$, which is therefore an optimal reconstruction distribution.